### v2

# Interconnect-Aware Logic Resynthesis for Multi-Die FPGAs

Abstract.
Multi-die FPGAs enable device scaling beyond reticle limits but introduce severe interconnect overhead across die boundaries. Inter-die connections, commonly referred to as super-long lines (SLLs), incur high delay and consume scarce interposer interconnect resources, often dominating critical paths and complicating physical design. To address this, this work proposes an interconnect-aware logic resynthesis method that restructures the LUT-level netlist to reduce the number of SLLs. The resynthesis engine uses die partitioning information to apply logic resubstitutions, which simplifies local circuit structures and eliminates SLLs. By reducing the number of SLLs early in the design flow, prior to physical implementation, the proposed method shortens critical paths, alleviates pressure on scarce interposer interconnect resources, and improves overall physical design flexibility. We further build a tool flow for multi-die FPGAs by integrating the proposed resynthesis method with packing and placement. Experimental results on the EPFL benchmarks show that, compared with a state-of-the-art framework, the proposed method reduces the number of SLLs by up to 24.8% for a 2-die FPGA and up to 27.38% for a 3-die FPGA. On MCNC benchmarks, our tool flow achieves an average SLL reduction of 1.65% while preserving placement quality. On Koios benchmarks, where fewer removable SLLs exist, several designs still exhibit considerable inter-die edge reductions. Overall, the results confirm that reducing inter-die connections at the logic level is an effective approach for multi-die FPGAs.
Interconnect, Logic Resynthesis, Multi-die FPGAs
$c$s: CCS Conceptsccs: Hardware Combinational synthesisccs: Hardware Partitioning and floorplanningccs: Hardware Physical synthesis
1. Introduction
Interposer-based multi-die integration has emerged as a practical solution to overcome the reticle size limits of lithography, enabling large systems to be implemented by partitioning designs across multiple dies (radojcic2017more, ). This approach is now widely adopted in both application-specific integrated circuits (ASICs) and field-programmable gate arrays (FPGAs(graening2025chipletpartcostawarepartitioning25d, ; Hahn2014fpga, ). However, distributing a design across dies inevitably introduces inter-die connections, i.e., links that cross boundaries between different dies, which are also referred to as super-long lines (SLLs).
Compared to intra-die connections that remain within a single die, inter-die connections raise two major challenges. First, they incur significantly larger delay because the signals need to traverse long physical distances across dies through the interposer and pass through additional interface circuitry. Thus, timing paths including inter-die connections are much slower, and even a small number of such connections can noticeably increase the critical path delay This effect is particularly profound in multi-die FPGAs, where signals can cross die boundaries only through dedicated inter-die links, making timing closure more difficult to achieve (boutros2023into, ; Nasiri2016TVLSI, ; Hahn2014fpga, ). Second, excessive inter-die connections increase physical design complexity. Inter-die connections consume scarce die-to-die links, which constrain block placement and tighten routing constraints. Prior work on single-die FPGAs has shown that higher interconnect complexity leads to longer physical design runtime (wang2024slip, ). In multi-die CAD flows, this effect can be further amplified. Inter-die connections tightly couple implementations on different dies, so physical implementation decisions on one die affect others. This coupling limits independent processing of dies, reduces parallelism in the design flow, and increases overall implementation difficulty and runtime.
To address these issues, prior work has explored reducing inter-die connections for FPGAs at different stages of the CAD flow. These efforts include multi-die-aware placement (LEAPS2024TCASI, ) and co-optimization of high-level synthesis directives with floorplanning (guo2021autobridge_fccm, ; FACO2023FPGA, ), highlighting that inter-die delay is a critical concern across different abstraction levels. Among them, partitioning-based FPGA CAD flows reduce inter-die connections through timing-aware assignment of blocks across dies (kuo2018pin, ; liao2020pin, ). Previous work integrates timing-driven partitioning into the open-source Verilog-to-Routing (VTR) framework (jasonluu2017vtr7, ), demonstrating improvements in critical-path delay and maximum operating frequency by discouraging cuts on timing-critical paths (iyer_partitioning-based_2025, ; raveena2024HEART, ).
While these methods can effectively reduce the inter-die connections, they operate on a netlist that has already been fixed by logic synthesis and technology mapping, meaning the underlying logic structure that generates the inter-die connections remains unchanged. Thus, they can only redistribute existing connections across dies, rather than re-evaluating whether inter-die connections are logically necessary. Even with high-quality partitioning, many inter-die connections may persist because partitioning does not change the circuit logic, so signals required on multiple dies must still be communicated across die boundaries. As a result, existing flows across different abstraction levels mitigate inter-die connections, rather than fundamentally reducing reliance on them from the logic level.
To overcome this limitation, we explore logic resynthesis as a means to directly modify the netlist connectivity at the logic level. Operating on mapped netlists, such as the lookup table (LUT)-level netlists commonly used in FPGA design, logic resynthesis modifies the netlist so that internal signals are computed and shared differently, using functional equivalences and don’t-care conditions. Conventionally, logic resynthesis has been used to optimize area, delay, or power, while we propose to leverage it to optimize inter-die connectivity in multi-die FPGAs. This logic-level restructuring capability introduces a new optimization dimension for reducing inter-die connections at their logical origin by selectively resynthesizing Boolean logic, thereby lowering SLL utilization and mitigating interposer-induced timing degradation.
The feasibility of this approach is supported by our benchmark analysis. We observe that a large fraction of signal connectivity originates from LUT-level soft logic rather than fixed architectural hard blocks. This indicates that many inter-die connections arise from flexible logic structures that can be modified, rather than rigid block-level interfaces. Consequently, even after partitioning, a large fraction of inter-die connections remains amenable to modification through logic-level restructuring.
Building on the above insight, we propose an interconnect-aware logic resynthesis method for multi-die FPGAs. Unlike existing approaches that operate at the partitioning or physical design stages, the proposed method directly eliminates unnecessary inter-die connections at their logical origin in the LUT-level netlists. To the best of our knowledge, this is the first work that optimizes inter-die connections for multi-die FPGAs at the logic resynthesis stage. Our main contributions are summarized as follows:
  1. (1)
    We formulate interconnect-aware logic resynthesis as an optimization problem that directly targets the reduction of SLLs in multi-die FPGAs.
  2. (2)
    To solve the optimization problem, we propose an SLL-aware resynthesis technique called ResynMD. It iteratively reconstructs local logic in a LUT-level netlist to eliminate unnecessary inter-die connections, while preserving functional correctness and avoiding area increase.
  3. (3)
    We develop a multi-die FPGA CAD tool flow that integrates the proposed resynthesis technique after partitioning and before physical design, which complements existing flows.
Experimental results show that on EPFL benchmarks, compared to a state-of-the-art partitioning-based multi-die FPGA CAD flow, our method reduces the number of SLLs by up to 24.8% for 2-die and 27.38% for 3-die FPGAs. On MCNC benchmarks, our method achieves an average reduction of 3.24% in inter-die connections, while preserving overall placement quality. On Koios benchmarks, the proposed method can also reduce inter-die connections while preserving post-placement quality.
The remainder of this paper is organized as follows. Section 2 describes the preliminaries. Section 3 elaborates the proposed interconnect-aware logic resynthesis method for multi-die FPGAs. The experimental results are presented in Section 4, followed by the conclusion and future work in Section 5.
2. Preliminaries
This section reviews the architectural and algorithmic background on interconnect-aware logic resynthesis for multi-die FPGAs. We first describe the characteristics of multi-die FPGA fabrics and the resulting non-uniform interconnect landscape, highlighting the role of inter-die connections and their impact on timing and physical design. We then summarize partitioning-based multi-die FPGA CAD flows. Finally, we introduce the LUT-level network representation.
Figure 1. Lateral and top views of a four-die FPGA system (AMD Xilinx Alveo U250 (amd-u200-ds962, )).
2.1. Multi-Die FPGAs
Multi-die FPGAs integrate multiple FPGA dies within a single package, typically using a silicon interposer to provide inter-die communication. This 2.5D integration paradigm enables device scaling beyond reticle limits, but results in non-uniform interconnect characteristics between intra-die and inter-die routing resources. Signals that cross die boundaries traverse interposer links, commonly modeled as SLLs, as illustrated in Figure 1.
Figure 2. Fabric-level interconnect abstractions for single-die and multi-die FPGAs: (a) a conventional single-die fabric; (b) a multi-die fabric with routing-based interposer interfaces; and (c) a multi-die fabric with tile-based interposer interfaces.
At the fabric level, this non-uniformity is reflected in the top-view architecture. As shown in Figure 2, a single-die FPGA provides a largely homogeneous routing fabric. In contrast, the multi-die fabric in Figure 2 exposes inter-die communication through a limited number of fixed-location, fixed-length SLLs connected to the routing fabric via dedicated switch boxes. Compared to intra-die routing wires, SLLs offer reduced routing flexibility and distinct delay characteristics.
Commercial multi-die FPGAs may further employ tile-based interposer interfaces with dedicated interface tiles (e.g., Laguna columns) (ravishankar2018sat, ) shown in Figure 2. Since the logic resynthesis techniques studied in this work operate on a generic logic netlist representation, they are agnostic to the specific implementation of the interposer. For clarity and compatibility with open-source academic CAD tools, we adopt the routing-based abstraction throughout this paper.
Table 1 summarizes representative delay values for intra-die routing wires and inter-die interposer links used in this work. The percentage indicates the fraction of routing tracks of each wire type within the total routing channel width. These values are derived from published technology-scaled delay models (Oleg2016Delay, ; Sun2019Pattern, ) and serve as parameters for characterizing the timing differences between routing resources.
Table 1. Intra-die (22 nm) and inter-die (45 nm interposer) wire delays.
Wire Type Technology Delay (ps) Percentage (%)
L1 22 nm (Intra-die) 76.2 50
L2 22 nm (Intra-die) 94.9 25
L6 22 nm (Intra-die) 212.0 25
L36 45 nm (Inter-die) 2223.7 48
2.2. Partitioning-Based Multi-Die FPGA Flow
By decomposing a monolithic design into multiple sub-circuits on multiple dies, partitioning reduces the optimization space of the CAD problems and enables scalable and partially parallel CAD processing, while still requiring careful coordination of inter-die connections. Figure 3 illustrates two representative partitioning-based CAD strategies for multi-die FPGA implementation, which differ in whether partitioning is performed before or after packing. In packing-first flows, LUTs are first clustered into architecture-specific blocks, such as configurable logic blocks (CLBs), and partitioning is subsequently applied to these packed blocks (iyer_partitioning-based_2025, ; graening2025chipletpartcostawarepartitioning25d, ). This ordering closely reflects physical constraints and simplifies downstream placement and routing, but fixes the logic connectivity at the granularity of packed blocks. In contrast, partitioning-first flows partition the primitive netlist before packing, allowing logic connectivity to be considered before architectural constraints are fully imposed. Several state-of-art multi-die CAD flows (iyer_partitioning-based_2025, ; graening2025chipletpartcostawarepartitioning25d, ; raveena2022slip, ; raveena2024HEART, ) use the partitioning-first paradigm. The netlist is first partitioned into sub-circuits on multiple dies, after which packing, placement, and routing are performed for each die.
Our interconnect-aware resynthesis technique is better suited to the partition-first flow. In this flow, we insert logic resynthesis after partitioning and before packing. Operating on the LUT-level netlist at this stage preserves more freedom for modification because LUTs have not yet been grouped into CLBs. Compared with a CLB-level netlist, the LUT-level representation is finer-grained and exposes more connections, allowing resynthesis to restructure logic and reduce inter-die connections more effectively.
(a) Packing-first flow.
(b) Partitioning-first flow.
Figure 3. Two partitioning-based CAD strategies for multi-die FPGAs: (a) a packing-first flow, where packing precedes partitioning; and (b) a partitioning-first flow, where the partitioning is performed before packing.
2.3. LUT-Level Network Representation
FPGA logic synthesis typically operates on circuits represented as graph-based logic networks, which provide a structural abstraction of Boolean functionality. After optimizing the logic networks, technology mapping transforms the circuit into a netlist of $k$-input LUTs. The LUT-level netlist explicitly defines the logic connectivity exposed to physical design processes, and therefore serves as the interface between logic synthesis and architecture-aware optimization. A LUT-level netlist can be viewed as a directed acyclic graph $G=(V,E)$. In this graph, each node $v\in V$ represents a $k$-input LUT implementing a local Boolean function $f_{v}$. Each directed edge $(u,v)\in E$ denotes a signal connection from the $u$ to $v$. Operating at the LUT level enables fine-grained optimization of inter-die connectivity by directly manipulating inter-LUT signals while ensuring seamless integration with downstream CAD flows.
For a node $v$, its transitive fanin (TFI) is the set of nodes that can reach $v$ through directed paths in $G$. Similarly, its transitive fanout (TFO) is the set of nodes reachable from $v$. These structural regions characterize the upstream and downstream influence of $v$ in the logic network. The maximum fanout-free cone (MFFC) of node $v$, denoted $\mathrm{MFFC}(v)$, is the maximal subgraph rooted at $v$ such that every node in the subgraph has no fanout outside the subgraph. Removing $v$ will disconnect all nodes in $\mathrm{MFFC}(v)$ from the primary outputs. The MFFC corresponds to the portion of logic exclusively controlled by $v$, and removing $v$ eliminates its entire MFFC.
3. Interconnect-Aware Resynthesis for Multi-Die FPGAs
This section presents ResynMD, an interconnect-aware logic resynthesis method for multi-die FPGAs. We first introduce the motivation and overall tool flow in Section 3.1. Then, Section 3.2 formulates the SLL-aware resynthesis problem and optimization objective. Section 3.3 describes the core components of the proposed resynthesis engine. Finally, Section 3.4 discusses the trade-off between partition imbalance and inter-die connection reduction.
3.1. Motivation and Overview
Modern applications mapped onto multi-die FPGAs typically contain large amounts of soft logic (i.e., logic implemented using LUTs) with many interconnections. Compared with hard blocks such as DSPs and RAMs, soft logic is highly flexible in placement and routing and is therefore more likely to be distributed across dies. While hard blocks generally reside at fixed sites and are rarely split across dies, LUT-level logic in large designs is frequently partitioned, producing numerous cross-die connections implemented using SLLs. Because SLLs incur high delay and limited routing flexibility, they frequently dominate critical paths and contribute significantly to routing congestion.
This work focuses on reducing the number of SLLs by targeting the logic-level origins of inter-die connectivity. These inter-die connections are not purely required by architectural constraints, but often stem from the logic implementation of the mapped LUT-level netlist. Consequently, modifying the logic implementation can often replace cross-die signals with functionally equivalent in-die implementations, thereby reducing the inter-die connections. By directly restructuring the logic to reduce inter-die connections, ResynMD tool flow complements existing partitioning-first CAD flows for multi-die FPGAs (raveena2024HEART, ; LEAPS2024TCASI, ) (see Figure 3(b)), which optimize inter-die connections through partitioning or placement without altering the circuit at the logic level.
Figure 4. The proposed ResynMD tool flow. Compared with the baseline partitioning-first CAD flows (Fig. 3(b)), this tool flow adds an SLL-aware resynthesis step between partitioning and physical implementation.
Figure 4 shows the ResynMD tool flow. It starts from a synthesizable HDL description and outputs a physical implementation for a multi-die FPGA. It first performs conventional logic synthesis and technology mapping to obtain a LUT-level netlist. The netlist is then partitioned across multiple dies using a timing-aware partitioner, which reduces inter-die connections on critical paths while balancing logic resources across dies. After partitioning, each LUT is assigned a fixed die label, resulting in a partition-annotated netlist in which inter-die connections correspond to nets spanning multiple dies. ResynMD, which is an SLL-aware logic resynthesis step, is then applied to this partition-annotated netlist. Its objective is to reduce the number of SLLs by restructuring netlist logic while preserving the fixed die assignments. Following resynthesis, the optimized netlist is split into per-die sub-netlists. Thereafter, the standard physical design stages, including packing, placement, and routing, are executed for each die.
Compared to the existing partitioning-first CAD flows, such as LiquidMD (raveena2024HEART, ; Dries2017LiquidFPT, ; Dries2018MultiPart, ), the key distinction of the proposed tool flow in Figure 4 is the SLL-aware logic resynthesis step that is applied after partitioning but before physical implementation, as illustrated by the blue segment in Figure 4. To the best of our knowledge, ResynMD is the first approach that explicitly targets logic-level restructuring for reducing inter-die connections. This introduces a new optimization dimension that is complementary to existing partitioning-based and placement-based multi-die optimization techniques and can provide synergistic benefits when integrated into current multi-die FPGA design flows. The following subsection details the problem formulation and the core components of the proposed SLL-aware resynthesis method.
3.2. Problem Formulation of SLL-Aware Logic Resynthesis
Conventional logic synthesis methods, such as  (mishchenko2006scalable, ; lee2021simulation, ; amaru2018improvements, ; liu2024cbtune, ), optimize metrics such as area or delay without considering the multi-die context and do not explicitly reduce inter-die connections. In contrast, the proposed SLL-aware logic resynthesis operates on a partitioned LUT-level netlist and directly targets inter-die connections. In this setting, the die assignment of each LUT is fixed by the partitioning stage, and only functionally equivalent logic transformations are permitted.
Let $G=(V,E)$ denote the LUT-level netlist with partition labels, where each node $v\in V$ is assigned to a die $D(v)$. A connection between nodes $u$ and $v$ is inter-die (and thus implemented using an SLL) if $D(u)\neq D(v)$. We define the total number of SLLs (also called inter-die connectivity) as
(1) $N_{sll}(G)=\sum_{(u,v)\in E}\mathbb{I}[\mathcal{D}(u)\neq\mathcal{D}(v)],$
where $\mathbb{I}[\cdot]$ is an indicator function that equals 1 if the connection between $u$ and $v$ crosses the die boundary and 0 otherwise.
Our objective is to construct a functionally equivalent netlist $G^{\prime}\equiv G$ that minimizes the number of SLLs without increasing the netlist area. Formally, the objective is
(2) $\min_{\begin{subarray}{c}G^{\prime}\equiv G\end{subarray}}N_{sll}(G^{\prime})\text{ subject to }Area(G^{\prime})\leq Area(G),$
where the function $Area(\cdot)$ measures the netlist area in terms of LUT count.
Because global optimizing Equation (2) over all equivalent netlists needs exploring a huge combinatorial search space, we adopt a greedy heuristic that iteratively applies local resubstitution (mishchenko2006scalable, ) transformations, which preserve functionality while modifying the fanin structure of nodes. This approach offers good scalability and computational efficiency for large-scale circuits. In each iteration, a transformation is accepted only if it strictly reduces the total number of SLLs and satisfies the area constraint. More details on the resubstitution procedure are presented in Section 3.3.
3.3. Details of SLL-Aware Logic Resynthsis
3.3.1. Greedy Algorithm for SLL-Aware LUT-Level Resynthesis
(a) Conventional MFS-based resubstitution.
(b) MFS-based interconnect-aware resubstitution.
Figure 5. Comparison between conventional MFS-based LUT-level resubstitution and the proposed interconnect-aware resubstitution.
Input: LUT-level netlist $G=(V,E)$, die mapping $\mathcal{D}(v)$ for each node $v\in V$
Output: Optimized LUT-level netlist $G^{\prime}$ with reduced number of SLLs.
1
2Initialize working netlist $G^{\prime}\leftarrow G$;
3
4for each pivot node $v$ in $G^{\prime}$ in topological order do
5  if $\forall u\in\text{fanin}(v):\mathcal{D}(u)=\mathcal{D}(v)$ then
6     continue; // Skip the node without cross-die fanins
7
8 Sub-circuit $W(v)\leftarrow\textit{BuildWindow(G', v)}$; // Extract local circuit surrounding $v$
9 Divisor set $\mathcal{S}_{d}(v)\leftarrow\textit{GetSLLAwareDivisors}(W(v),v)$; // Get divisors used to re-express $v$
10 New function $\hat{f}_{v}\leftarrow\textit{FindEquivFunc}(W(v),v,\mathcal{S}_{d}(v))$; // Find $v$’s equivalent function
11 if $\hat{f}_{v}$ exists then
12     New node $v^{\prime}\leftarrow LUTImplement(\hat{f}_{v})$;
13
14    if #\#cross-die fanins of $v^{\prime}<\#$cross-die fanins of $v$
15     and $Area(v^{\prime})\leq Area(v)$ then
16
17       $G^{\prime}\leftarrow\textit{ApplyResubstitution}(G^{\prime},v,v^{\prime})$; // Replace $v$ with new node $v^{\prime}$
18
19
20return $G^{\prime}$;
Algorithm 1 Greedy SLL-Aware Logic Resynthesis by LUT-level Resubstitution
This subsection presents the core algorithm of the proposed SLL-aware logic resynthesis method. The method builds upon the resynthesis technique MFS (alanmi2011mfs, ). Unlike conventional MFS, which optimizes traditional metrics such as area or delay, our new contributions lie in adapting the resynthesis approach to the multi-die context with a specific focus on reducing inter-die connections, as illustrated in Figure 5. The inter-die connections are reduced by performing LUT-level resubstitution. Here, resubstitution (alanmi2011mfs, ; mishchenko2006using, ) refers to a local logic transformation that replaces a node with a functionally equivalent implementation expressed in terms of a selected set of other nodes (called divisors) in the netlist.
As shown in Algorithm 1, the input of our resynthesis procedure is a partition-annotated LUT-level netlist $G$, where each node is assigned a die label determined during partitioning. The output is an optimized netlist $G^{\prime}$ with fewer inter-die connections, which can be directly fed into the subsequent physical implementation stages. The algorithm initializes the working netlist $G^{\prime}$ as the input netlist $G$ (Line 1). Then, it processes each node $v$ in $G^{\prime}$ in topological order (Line 1), ensuring that fanin nodes are handled before their fanouts and preventing cyclic dependencies during resubstitution. We call the node $v$ being processed the pivot node. For each pivot node, the algorithm first checks whether all fanins of $v$ reside on the same die as $v$ (Line 1). If so, node $v$ has no cross-die fanins and cannot contribute to SLL reduction, and is therefore skipped (Line 1). For node $v$ with cross-die fanins, a local window $W(v)$ is constructed (Line 1). This window is a sub-circuit surrounding node $v$ that captures the local logic context of $v$ while limiting the scope of optimization to ensure computational tractability. Within this window, a set of SLL-aware divisors $\mathcal{S}_{d}(v)$ is collected (Line 1), which are potential candidate signals used to construct $v$’s new implementation. These divisors consist only of nodes located on the same die as $v$, ensuring that any new implementation of $v$ can be realized without introducing new cross-die fanins and enabling the elimination of existing ones. Using these divisors, the algorithm attempts to derive an equivalent Boolean function $\hat{f}_{v}$ for node $v$ (Line 1). Note that such a function may not always exist (see Section 3.3.4 for details). If no such function exists, the node $v$ remains unchanged. Otherwise, the function $\hat{f}_{v}$ is implemented using one or more LUTs to create a new node $v^{\prime}$ (Line 1), which serves as a candidate resubstitution for $v$. The candidate resubstitution is accepted only if the new node $v^{\prime}$ has fewer cross-die fanins than the original node $v$, and does not increase area in terms of LUT count (Lines 1 and 1). When both conditions are satisfied, the resubstitution is committed to the working netlist $G^{\prime}$ (Line 1), and the algorithm proceeds to the next node. After all nodes have been processed, the optimized netlist $G^{\prime}$ is returned (Line 1).
The above greedy strategy guarantees monotonic reduction of cross-die fanins of each node $v$ in the LUT-level netlist while maintaining scalability for large designs. Since each accepted resubstitution removes at least one inter-die fanin edge of node $v$ (Line 1), the total number of SLLs $N_{sll}$ decreases monotonically ensuring the convergence of the procedure. Moreover, the area constraint in Equation (2) is enforced through the condition check in Line 1, which disallows any resubstitution that increases LUT count.
Algorithm 1 relies on three key procedures corresponding to three main components of the SLL-aware resynthesis method: BuildWindow, which constructs the local optimization region, GetSLLAwareDivisors, which selects same-die candidate signals, and FindEquivFunc, which constructs an equivalent function, typically via SAT-based resubstitution. These components are described in detail in Section 3.3.23.3.4.
3.3.2. Window Construction
This step corresponds to Line 1 in Algorithm 1. We adopt the same window construction strategy as MFS (alanmi2011mfs, ). As illustrated in the middle of Figure 5(b), for each pivot node $v$ (i.e., node being processed for resubstitution), its window $W(v)$ is a sub-circuit surrounding $v$ that captures the local logic context relevant for $v$’s resubstitution. $W(v)$ is formed by expanding its TFOs up to a limited depth $D_{1}$. Then, $D_{2}$ levels of TFIs of the pivot node $v$ determines the window primary inputs (PIs). After that, the internal nodes are obtained through a topologically ordered traversal of the induced subgraph, forming the window $W(v)$.
3.3.3. Divisor Selection
This step corresponds to the computation of the divisor set $\mathcal{S}_{d}(v)$ in Line 1 of Algorithm 1, where candidate divisors are extracted from the local window $W(v)$ and filtered according to their die assignment. Divisors are signals within the local window $W(v)$ that can be used to re-express the Boolean function of the pivot node $v$. Their selection process is illustrated in the middle part of Figure 5(b). When cross-die fanins of the pivot are detected, a local window is constructed around the pivot node. Window PIs are split according to whether they belong to the transitive fanin of the pivot. The nodes on structural paths between the pivot and its TFI-related window PIs are collected as potential divisors, excluding the pivot itself and the nodes in its maximum fanout-free cone (MFFC). Additionally, window nodes whose structural support does not depend on non-TFI window PIs are also considered. To limit complexity, divisors that exceed a predefined logic-level bound are discarded, and the total number of divisors is capped.
To incorporate inter-die interconnection awareness, each divisor is classified according to its die assignment relative to the pivot node. The divisors located on the same die (highlighted in dark green in Figure 5(b)) as the target are greedily selected and constitute the final in-die divisor set $S_{d}(v)$.
In conventional resubstitution, candidate divisors are selected solely based on functional or structural criteria without considering their physical die assignment, sometimes introducing new inter-die connections. In contrast, the proposed approach enforces a greedy, SLL-aware divisor selection strategy that deterministically considers same-die signals. This enables the following SAT-based resubstitution approach to eliminate unnecessary inter-die connections.
3.3.4. SAT-Based Resubstitution
This step corresponds to the Line 1 of Algorithm 1. The goal is to compute a function $\hat{f}$ that is equivalent to $f_{v}$ (the function of the pivot node $v$), expressed using a subset of in-die divisors $S_{d}^{\prime}\subseteq S_{d}(v)$. The key idea is to selectively eliminate cross-die fanins of the pivot node $v$ by expressing its function $f_{v}$ using only in-die signals. For a pivot node $v$, we construct its equivalent function $\hat{f}$ in a structured manner, as summarized in Algorithm 2. Rather than searching for arbitrary divisor combinations, we restrict the construction to eliminate cross-die dependencies.
Input: Local window $W(v)$, pivot node $v$, divisor set $\mathcal{S}_{d}(v)$
Output: Equivalent function $\hat{f}$ or $\emptyset$ (i.e., $\hat{f}$ does not exist)
1
2Select a cross-die fanin $u\in\text{fanin}(v)$;
3
[…]
5
[…]
7if $\textit{SATBasedExistCheck}(f_{v},C,S_{d}^{\prime})$ then
8  $S_{d}^{\prime}\leftarrow BaseDivisors$;
9  return $\hat{f}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\leftarrow}}\textit{Interpolate}(S_{d}^{\prime})$;
10
11
12for each divisor $d\in\mathcal{S}_{d}(v)$ do
13  $S_{d}^{\prime}\leftarrow BaseDivisors\cup\{d\}$;
14  if $\textit{SATBasedExistCheck}(f_{v},C,S_{d}^{\prime})$ then
15     return $\hat{f}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\leftarrow}}\textit{Interpolate}(S_{d}^{\prime})$;
16
17
18
19return $\emptyset$;
20
[…]
First, this procedure is triggered by identifying a cross-die fanin $u\in\mathrm{fanin}(v)$ and attempting to remove it. The base divisor subset $S_{d}^{\prime}=BaseDivisors$ is initialized by excluding $u$ while retaining all remaining fanins (Line 2). If $f_{v}$ can be expressed using this reduced set under the care constraint, the cross-die edge $(u,v)$ is directly eliminated.
If this direct elimination fails, we traverse the in-die divisor set $S_{d}(v)$ (Line 2). For each candidate divisor $d$, we construct a new divisor subset $S_{d}^{\prime}$ by augmenting the base divisors with $d$, i.e., by adding a single in-die divisor to compensate for the removed cross-die dependency. Feasibility is then re-evaluated for this augmented set.
For each constructed divisors subset $S_{d}^{\prime}$, a SAT-based existence check (alanmi2011mfs, ) is performed under the extracted care set $C$ to determine whether $f_{v}$ can be equivalently expressed using signals in $S_{d}^{\prime}$. If the SAT-based existence check succeeds, an interpolation is extracted to construct the equivalent function $\hat{f}_{v}$ (lee2007interpolation, ). The interpolation depends only on signals in $S_{d}^{\prime}$ and preserves equivalence under the care set $C$. The derived function $\hat{f}_{v}$ is then implemented as a new LUT node $v^{\prime}$ whose fanins correspond to the selected divisors in $S_{d}^{\prime}$. Node $v^{\prime}$ is inserted into the original network, and all fanout connections of the original pivot node $v$ are redirected to $v^{\prime}$. Subsequently, node $v$ and any newly unreachable logic in its $\mathrm{MFFC}(v)$ are removed from the network. This network update corresponds to line 1 in Algorithm 1.
3.3.5. Example
We use a small 2-input LUT (LUT2 in short) netlist as an example to illustrate how our SLL-aware resynthesis technique can reduce redundant inter-die connections. Note that LUT2 is used here for ease of presentation, and the proposed method is applicable to general LUT sizes.
[…]
The logic functions in the 2-input LUTs are defined as:
(3) $X$ $=a\oplus b$
$Y$ $=X\oplus c$
$F$ $=a\oplus d$
(a) Pre-resynthesis: the original LUT-level netlist before resubstitution ($F$ is pivot node).
(b) After resynthesis: SLL-aware resubstitution eliminating the redundant inter-die dependency on $a$.
Figure 6. Example of SLL-aware resynthesis.
Don’t-care and care set definition. As discussed in Section 3.3.4, the functionality of a pivot node is required to be preserved only on its care set, and the don’t-care cases can be exploited to enable more flexible resubstitution. For illustrative purposes, we assume that input assignments with $b\neq c$ are don’t-cares of the pivot node $F$. In other words, we only care about the functional equivalence under the input cases when $b=c$. The don’t-care and care conditions are defined as:
(4) $DC(a,b,c,d)$ $=b\oplus c,$
$C(a,b,c,d)$ $=\lnot(b\oplus c).$
Resynthesis process and post-resynthesis logic. Node $F$ is the pivot node to start the process. As the current example is simple, following the windowing method introduced in Section 3.3.2, all of the LUT nodes across two dies will be covered in the window. During the divisor selection step, within the resynthesis window, only the in-die signals are available as candidate divisors, including $Y$, $c$ and $d$. Exploiting the don’t-care conditions in Equation (4) and the candidate divisors, we can use the method in Section 3.3.4 to construct an equivalent function of $F$. First, the cross-die fanin $a$ is selected and excluded, while the remaining fanin $d$ is extracted as $BaseDivisors$. Following Algorithm 2, two types of divisor subsets are examined. The first consists only of $BaseDivisors$ (i.e., $S_{d}^{\prime}=\{d\}$). The second augments $BaseDivisors$ with an additional in-die divisor $Y$, resulting in $S_{d}^{\prime}=\{d,Y\}$. The SAT-based existence check returns true for $S_{d}^{\prime}=\{d,Y\}$, indicating that $F$ can be equivalently expressed using only in-die divisors. The interpolated function is therefore derived as
(5) $F^{\prime}=Y\oplus d.$
As shown in Table 2, the new function $F^{\prime}$ is equivalent to $F$ on all care minterms, which ensures the functional correctness after resubstitution. After resubstituting $F$ by $F^{\prime}$, the post-resynthesis netlist is shown in Figure 6(b). $F^{\prime}$ no longer depends on the inter-die signal $a$, while the necessary inter-die dependency on $X$ is preserved via node $Y$. The amount of SLLs $(N_{sll})$ of this small example decreases from 2 to 1.
Table 2. Truth table of the pivot node before and after resubstitution. Gray rows correspond to don’t-care conditions ($b\neq c$). Signals $Y$ and $d$ are in-die divisors used to reconstruct $F$.
$a$ $b$ $c$ $d$ $X$ $Y$ $F$ (orig.) $F^{\prime}$ (new) care?
0 0 0 0 0 0 0 0 1
0 0 0 1 0 0 1 1 1
0 0 1 0 0 1 0 1 0
0 0 1 1 0 1 1 0 0
0 1 0 0 1 1 0 1 0
0 1 0 1 1 1 1 0 0
0 1 1 0 1 0 0 0 1
0 1 1 1 1 0 1 1 1
1 0 0 0 1 1 1 1 1
1 0 0 1 1 1 0 0 1
1 0 1 0 1 0 1 0 0
1 0 1 1 1 0 0 1 0
1 1 0 0 0 0 1 0 0
1 1 0 1 0 0 0 1 0
1 1 1 0 0 1 1 1 1
1 1 1 1 0 1 0 0 1
3.4. Partition Imbalance Induced by Inter-Die Resynthesis
In addition to minimizing the number of inter-die connections, partitioning-based multi-die FPGA CAD flows are also expected to maintain a balanced distribution of logic resources across dies. Partition balance is a key feasibility constraint: a severely imbalanced die can become overutilized, leading to packing/placement failures or degraded QoR. A commonly used metric is the partition imbalance ratio, denoted by $\rho$, which quantifies the deviation of logic distribution from an ideal uniform partition (raveena2022slip, ). It is defined as
(6) $\rho=\max_{k\in\{0,\dots,K-1\}}\left(\frac{|V_{k}|}{|V|/K}\right),$
where $K$ is the number of dies, $|V_{k}|$ denotes the number of logic nodes assigned to die $k$, and $|V|$ denotes the total number of logic nodes. In practice, an upper bound on $\rho$ called $UB$ is imposed during partitioning to ensure feasible packing and placement on each die.
Note that the proposed SLL-aware resynthesis does not strictly preserve the original partition balance. The success of resubstitution is inherently opportunistic and each success has an asymmetric and localized impact on the involved dies. Figure 5(a) illustrates a representative scenario. On the destination die, the pivot node $v$ is replaced by another node $v^{\prime}$, so the size typically remains unchanged. On the source die, the effect depends on the fanout structure. If the predecessor $v$ has no other fanouts, both $v$ and its MFFC become redundant and can be eliminated, thereby reducing the size of the corresponding die by one or more nodes. Otherwise, no node reduction occurs. This size reduction behaviour is the original purpose of conventional MFS and is inherited by SLL-aware resynthesis, which consequently introduces uncertainty in die imbalance. As a result, the number of logic nodes assigned to individual dies may change, leading to a variation in the observed imbalance factor $\rho$. Since resubstitution is applied greedily in topological order, such node removals are difficult to predict analytically. Although the total number of LUTs in the netlist is guaranteed to be non-increasing (Line 1 in Algorithm 1), the per-die logic distribution may vary slightly. Therefore, we report the imbalance factor before and after resynthesis in Sections 4.3 and 4.5 to evaluate its practical impact.
4. Experimental Results
This section evaluates the effectiveness of the proposed interconnect-aware resynthesis method, ResynMD, from both logic-level and physical design perspectives. We first assess the capability of ResynMD to reduce inter-die connectivity on purely combinational benchmarks, isolating the impact of logic resynthesis without interference from physical design effects. We then study the sensitivity of the proposed approach to LUT size, which directly influences resubstitution opportunities and optimization effectiveness. Finally, we evaluate the downstream impact of interconnect-aware resynthesis on partitioning-based multi-die placement, including wirelength, load balance, and runtime, using realistic heterogeneous FPGA benchmarks.
4.1. Experimental Setup
Our interconnect-aware logic resynthesis method for multi-die FPGAs, ResynMD, is implemented in C++ and Java. All experiments are conducted on a workstation equipped with an AMD Ryzen 9 7900X 12-core processor and 128 GB RAM, running a 64-bit Linux system. For the logic synthesis and technology mapping stage in Figure 4, we use the Yosys (yosys2025yosys, ) and ABC (brayton2010abc, ) tools. For the partitioning stage, we utilize an hMETIS-based (karypis1998hypergraph, ) timing-aware partitioning tool. For the physical implementation stage, we utilize the AAPack (Luu2011AAPack, ) clustering algorithm provided in VTR8 (murray_vtr_2020, ), and the multi-die placement algorithm implemented in LiquidMD (raveena2024HEART, ; Dries2017LiquidFPT, ). No post-routing results are reported in this work, while our tool flow can be extended to evaluate routing performance in future work.
We fix the imbalance factor $UB$ of the partitioner to 1.25 across all experiments. We evaluate our approach using three benchmark suites, each targeting a distinct evaluation objective. The EPFL (epfl, ) benchmark suite consists of purely combinational designs and is used to quantify the intrinsic effectiveness of the proposed logic-level resynthesis in reducing inter-die connections. Since these circuits are relatively small and do not reflect realistic physical-design constraints, we report only connectivity-level metrics on EPFL benchmarks.
To evaluate physical-design impact in the presence of sequential logic, we additionally use a set of MCNC (yang1991logic, ) benchmarks. These circuits contain flip-flops but no heterogeneous hard blocks, which enables isolating the effect of sequential elements while still allowing post-placement evaluation.
To evaluate physical-design impact on realistic heterogeneous FPGA designs, we utilize the Koios (arora2021koios, ) benchmark suite, which comprises designs featuring hard blocks such as DSPs and block RAMs. In these circuits, hard macros and timing-optimized sequential elements are preserved and treated as structural boundaries, allowing us to assess whether SLL-aware resynthesis remains beneficial under constrained rewriting flexibility. For both MCNC and Koios, we report post-placement results to quantify downstream physical-design effects.
4.2. Multi-Die Cost Modeling and Connectivity Metrics
During placement, wirelength is commonly estimated by summing the contributions of all nets, where the wirelength of each net is approximated using the half-perimeter wirelength (HPWL) of its bounding box. To account for net-specific characteristics, the HPWL of each net is multiplied by a weighting factor $q(n)$ (Dries2017LiquidFPT, ; raveena2024HEART, ). The overall wirelength cost is thus defined as
(7) $BBox_{cost}^{SD}=\sum_{n\in\text{nets}}q(n)\cdot\mathrm{HPWL}(n).$
for a single-die FPGA.
Figure 7. Bounding-box ($B$) wirelength estimation before and after die partitioning (raveena2022slip, ). (a) A non-partitioned net enclosed by a single bounding box within one die. (b) After partitioning, the net spans multiple dies, forming die-local terminal clusters. (c) The partitioned net is modeled using multiple die-local bounding boxes connected by inter-die segments, and the total $B$ wirelength accounts for both intra-die and inter-die components.
For multi-die FPGA systems, Figure 7 illustrates how bounding-box wirelength estimation changes before and after die partitioning. For a non-partitioned design, all terminals of a net are enclosed by a single bounding box within one die (Figure 7(a)). After partitioning, where partitions equate to dies—nets are classified as intra-die (terminals within one die) or inter-die (crossing boundaries). Intra-die nets use standard HPWL within the die region. For inter-die nets, a single bounding box underestimates wirelength due to long interposer links. Thus, each net is decomposed into die-local segments (evaluated via HPWL with virtual boundary terminals) and an interposer segment modeled as a fixed-length SLL.
Accordingly, the total estimated wirelength of the multi-die system is computed as the sum of the per-die bounding-box costs defined in Equation (7), together with the contribution of inter-die connections, and is expressed as
(8) $\mathrm{BBox}_{\text{cost}}^{MD}=\sum_{d\in\mathcal{D}}BBox_{cost}^{d}+N_{sll}\cdot L_{sll},$
where $N_{sll}$ denotes the number of nets crossing the die boundary and $L_{\text{sll}}$ is the length of a single interposer link. This modular formulation enables independent optimization of each die during placement while accurately accounting for the cost of timing-expensive inter-die connections.
4.2.1. Multi-Pin Inter-Die Nets and Edge-Level Objective
For a multi-pin net whose sinks are distributed across different dies, current physical design flows typically introduce dummy blocks (e.g., SLL buffers) to model inter-die connections. These dummy blocks are placed at the source and sink sides of the SLLs, enabling consistent modeling in both tile-based and routing-based multi-die FPGA architectures. Furthermore, these cross-die sinks can share the same SLL channel, thereby reducing the second term of the system bounding-box cost in Equation 8.
Figure 8. Illustration of a multi-pin inter-die net implemented via SLL dummy blocks. Eliminating the dependency on node $n$ removes one inter-die connection (i.e., reduces the SLL net’s fanout by one), while the SLL net itself remains due to other cross-die sinks.
As illustrated in Figure 8, in this example, an inter-die net may fan out to multiple sinks on the destination die through an SLL modeled as a dummy block. Each successful resubstitution in resynthesis is pivot-centric: the optimization is performed with respect to a selected pivot node, and only its fanin structure is modified. Consequently,the resub removes at most the inter-die fanin edges incident to that pivot, while leaving the fanout structure of the source net unchanged. As illustrated in Figure 8, when node $n$ is selected as the pivot, resubstitution can eliminate its dependency on the inter-die signal, other sinks of the same net remain unaffected. Therefore, a single resubstitution reduces exactly one SLL net’s fanout ($SLL_{FO}$), but does not eliminate the entire SLL net if additional cross-die sinks still exist. Therefore, inter-die connectivity ($N_{sll}$) after physcial design implementation does not fully capture the structural impact of SLL-aware resynthesis. For this reason, we refine the definition of inter-die connectivity, using two metrics: the net-level count $N_{\text{sll}}$ and the edge (connection)-level SLL fanout count $N_{\text{sll}}^{FO}$.
The $N_{\text{sll}}$ counts the number of SLL nets occupying SLL tracks after physical implementation, where multiple cross-die sinks may share the same SLL channel. In contrast, $N_{\text{sll}}^{FO}$ counts each inter-die fanout edge in the LUT-level netlist individually. Each accepted resubstitution guarantees the removal of at least one inter-die edge, strictly reducing $N_{\text{sll}}^{FO}$ even if $N_{\text{sll}}$ remains unchanged.
In what follows, we evaluate the effectiveness of ResynMD on inter-die connectivity in Section 4.3 and its impact on physical design in Section 4.5.
4.3. Impact on Inter-Die Connectivity
We evaluate the effectiveness of ResynMD on pure combinational logic benchmarks from the EPFL benchmark suite. The input netlists are first partitioned using KaHyPar (KaHyPar, ) with equal node weights, without considering timing criticality. Subsequently, Algorithm 1 is applied to perform interconnect-aware logic resynthesis. As illustrated in Figure 9, the inter-die connectivity before and after interconnect-aware resynthesis highlights the effectiveness of the proposed optimization, where plenty of inter-die connections (shown in red) are eliminated.
Tables 3 and 4 summarize the impact of resynthesis on inter-die connectivity and partition balance under both 2-die and 3-die configurations. Across the evaluated benchmarks, ResynMD reduces the number of inter-die connections in a substantial fraction of cases (11/20 for 2-die and 12/20 for 3-die), while leaving others unchanged, which reflects the circuit-dependent nature of resubstitution opportunities.
(a) Inter-die connectivity.
(b) Partial optimized multi-pin inter-die net.
Figure 9. Inter-die connectivity of a two-die system for the voter benchmark from the EPFL suite. Logic blocks are grouped by die. Solid blue edges indicate inter-die connections that remain after resynthesis, while dashed red edges denote inter-die connections eliminated by interconnect-aware resynthesis.
Table 3. Inter-die connections ($N_{\text{sll}}$) and imbalance factor $\rho$ before and after ResynMD on EPFL benchmarks (2-die configuration).
Bench. $N_{\text{sll}}$ $\rho$
Pre Resyn $\Delta$ $\Delta(\%)$ Pre Resyn $\Delta$ $\Delta(\%)$
cavlc 42 41 -1 -2.38 1.0070 1.0277 +0.0207 +2.05
arbiter 5430 5171 -259 -4.77 1.0007 1.0000 -0.0007 -0.07
voter 57 32 -25 -43.86 1.0380 1.0010 -0.0370 -3.57
mem_ctrl 1136 815 -321 -28.26 1.0147 1.0161 +0.0014 +0.14
bar 526 458 -68 -12.93 1.0052 1.0084 +0.0032 +0.32
sin 321 248 -73 -22.74 1.0161 1.0650 +0.0489 +4.81
max 213 211 -2 -0.94 1.0374 1.0209 -0.0165 -1.59
square 57 39 -18 -31.58 1.0194 1.0354 +0.0160 +1.56
mult 1000 919 -81 -8.10 1.0003 1.0386 +0.0383 +3.83
log2 3032 1554 -1478 -48.74 1.0328 1.1083 +0.0755 +7.31
hyp 851 279 -572 -67.21 1.0252 1.0221 -0.0031 -0.30
int2float 11 11 0 0 1.0698 1.0698 0 0
Avg. -263.45 -24.80 +0.0133 +1.33
Table 4. Inter-die connections ($N_{\text{sll}}$) and imbalance factor $\rho$ before and after ResynMD on EPFL benchmarks (3-die configuration).
Bench. $N_{\text{sll}}$ $\rho$
Pre Resyn $\Delta$ $\Delta(\%)$ Pre Resyn $\Delta$ $\Delta(\%)$
cavlc 61 56 -5 -8.20 1.0385 1.0316 -0.0068 -0.66
arbiter 6322 6259 -63 -1.00 1.0042 1.0047 +0.0005 +0.05
voter 81 41 -40 -49.38 1.0435 1.1038 +0.0604 +5.79
mem_ctrl 2233 1807 -426 -19.09 1.0770 1.0800 +0.0030 +0.28
bar 816 671 -145 -17.77 1.0675 1.0967 +0.0293 +2.74
sin 1412 1123 -289 -20.48 1.0426 1.2183 +0.1758 +16.86
max 311 296 -15 -4.82 1.0187 1.0330 +0.0143 +1.40
square 78 60 -18 -23.08 1.0345 1.0241 -0.0104 -1.00
mult 1846 1704 -142 -7.69 1.0021 1.0721 +0.0700 +6.98
log2 6930 3146 -3784 -54.61 1.0299 1.0591 +0.0292 +2.83
hyp 1276 395 -881 -69.04 1.0253 1.0303 +0.0050 +0.49
int2float 20 19 -1 -5.00 1.1163 1.1163 0.0000 0.00
Avg. -467.33 -27.38 +0.0251 +2.35
[…]
The increased effectiveness is likely accompanied by a larger variation in partition imbalance. Compared to the 2-die case, the 3-die configuration exhibits a higher average increase in imbalance ratio calculated from Equation 6 ($2.35\%$ versus $1.33\%$), reflecting the more extensive structural changes introduced by resynthesis. Nevertheless, all observed imbalance values remain within practical bounds, and stricter imbalance constraints can be imposed during the initial partitioning stage if needed. Figure 10 further illustrates the relationship between inter-die connectivity reduction and imbalance variation. No strong correlation is observed between $\Delta\rho$ and $\Delta N_{sll}$, and several circuits achieve substantial SLL reduction with negligible imbalance change. This suggests that the effectiveness of ResynMD does not inherently rely on increasing partition imbalance.
[…]
Figure 10. Trade-off between inter-die connectivity reduction and imbalance variation under 2-die and 3-die configurations.
4.3.1. Physical Implications of Edge-Level SLL Reduction
Although ResynMD does not always reduce the number of SLL nets ($N_{\text{sll}}$), it consistently reduces the number of inter-die connections ($N_{\text{sll}}^{FO}$). This distinction is important from a physical design perspective. As illustrated by the green zoomed-in nets in Figure 9(b), one inter-die connection is removed, while another remains. High-fanout inter-die nets impose disproportionate routing and timing pressure, as they load inter-die resources and near-die boundary routing resources, and increase routing complexity within the destination die.
Reducing the number of inter-die SLL fanouts ($N_{\text{sll}}^{FO}$) decreases SLL loading and mitigates routing congestion around SLL endpoints. Since SLL sink locations are fixed at the die boundary and already incur substantial inter-die delay, signals traversing SLLs typically exhibit high arrival times and timing criticality. Under congestion, routing detours near these fixed interface points can further increase delay and exacerbate timing violations. By reducing inter-die fanout, the proposed method alleviates local congestion pressure around SLL interfaces, thereby improving routability and timing robustness even when the total number of occupied SLL tracks remains unchanged.
4.4. Impact of LUT Size on SLL-Aware Resynthesis Effectiveness
The effectiveness of SLL-aware resynthesis is strongly influenced by the target LUT size, which directly determines structural flexibility and divisor availability in the mapped LUT-level netlist. Smaller LUT sizes produce finer-grained logic decomposition, resulting in a larger number of nodes with simpler Boolean functions. This increases the number of candidate divisors within each resubstitution window and reduces SAT constraint complexity, thereby improving the likelihood of finding valid in-die replacements. In contrast, larger LUT sizes consolidate logic into fewer nodes with higher functional complexity. This reduces the available divisor set and enlarges the SAT search space, making it more difficult to identify feasible substitutions that reduce inter-die connectivity.
To quantify this sensitivity, we map the EPFL benchmarks to LUT-4/5/6 and measure the reduction in inter-die SLL nets after resubstitution, as summarized in Table 5. In this experiment, we use a larger window size than in the previous effectiveness study to avoid overly restricting the divisor space. Unlike the previous experiments that use min-cut partitioners (KaHyPar/hMetis), the initial die assignment here is generated by a deterministic hash-based labeling scheme to eliminate partitioning variability across different LUT mappings. Since this pseudo-partition is not optimized for cut minimization, the initial inter-die connectivity is typically higher, providing more opportunities for resubstitution to remove cross-die nets. As a result, the absolute SLL reductions are larger than those in Section 4.3, but still adequate for comparing LUT-size impact.
On average, LUT4 achieves the largest reduction (2243.42 nets), followed by LUT5 (1838.26), while LUT6 shows the smallest improvement (1622.11). This trend is consistent across most benchmarks, indicating that LUT granularity plays a critical role in interconnect-aware optimization. Post-partition resubstitution benefits from finer-grained logic representations, while larger LUT sizes limit optimization opportunities due to reduced structural flexibility.
[…]
Bench. LUT4 LUT5 LUT6 Bench. LUT4 LUT5 LUT6
int2float 46 33 25 sin 1382 1259 1155
ctrl 23 6 5 max 568 477 388
router 37 30 30 square 3789 3351 2949
cavlc 179 115 53 sqrt 3255 3205 2660
priority 145 113 105 multiplier 5107 4795 4328
dec 32 31 31 log2 6889 6141 5610
i2c 177 129 104 div 4538 3290 4032
arbiter 2215 1729 1417 voter 1960 1772 1425
mem_ctrl 9361 7374 6077
Total 42625 34927 30820 Average 2243.42 1838.26 1622.11
4.5. Impact on Physical Design (Placement)
[…]
Table 6. Impact of ResynMD on inter-die connectivity, placement quality, load balance, and placement runtime across MCNC benchmarks.
Bench. $N_{\text{sll}}$ $N_{\text{sll}}^{FO}$ $BBox_{cost}^{2die}$ $\rho_{pre}$ $\rho_{post}$ $\text{Runtime}_{place}$
P R $\Delta$ $\%$ P R $\Delta$ $\%$ P R $\Delta$ $\%$ P R $\Delta$ $\%$ P R $\Delta$ $\%$ P R $\Delta$ $\%$
apex2 66 64 -2 -3.03 400 293 -107 -26.75 582.9 575.3 -7.6 -1.30 1.130 1.129 -0.0005 -0.04 1.128 1.123 -0.005 -0.41 0.580 0.564 -0.016 -2.73
apex4 134 134 0 0.00 872 827 -45 -5.16 601.8 594.5 -7.3 -1.21 1.151 1.146 -0.005 -0.41 1.131 1.138 +0.007 +0.60 0.517 0.522 +0.005 +0.99
bigkey 6 6 0 0.00 784 784 0 0.00 459.9 424.8 -35.1 -7.64 1.006 1.006 0.000 0.00 1.000 1.000 0.000 0.00 0.644 0.651 +0.006 +1.01
clma 143 141 -2 -1.40 3821 2879 -942 -24.65 1996.4 2017.8 +21.4 +1.07 1.142 1.135 -0.007 -0.59 1.141 1.142 +0.002 +0.14 0.791 0.791 0.000 -0.01
des 11 11 0 0.00 581 581 0 0.00 325.0 346.3 +21.3 +6.56 1.026 1.026 0.000 0.00 1.025 1.024 -0.0002 -0.02 0.723 0.706 -0.018 -2.45
diffeq 49 49 0 0.00 995 980 -15 -1.51 468.0 465.8 -2.2 -0.47 1.130 1.130 0.000 0.00 1.128 1.134 +0.006 +0.53 0.545 0.539 -0.006 -1.17
dsip 3 3 0 0.00 560 560 0 0.00 281.0 275.5 -5.4 -1.94 1.008 1.008 0.000 0.00 1.006 1.006 0.000 0.00 0.624 0.627 +0.003 +0.52
ex5p 186 165 -21 -11.29 922 679 -243 -26.36 758.3 673.8 -84.5 -11.14 1.167 1.145 -0.022 -1.90 1.155 1.126 -0.029 -2.51 0.559 0.538 -0.021 -3.82
frisc 173 173 0 0.00 3075 2772 -303 -9.85 1142.8 1127.8 -15.0 -1.31 1.064 1.064 0.000 0.00 1.054 1.052 -0.002 -0.18 0.639 0.633 -0.006 -0.94
misex3 117 116 -1 -0.85 806 693 -113 -14.02 596.3 595.9 -0.4 -0.07 1.157 1.144 -0.013 -1.11 1.148 1.149 +0.002 +0.14 0.536 0.533 -0.003 -0.60
pdc 312 303 -9 -2.88 5577 4547 -1030 -18.47 1469.9 1411.7 -58.2 -3.96 1.171 1.155 -0.015 -1.31 1.152 1.186 +0.033 +2.89 0.644 0.661 +0.017 +2.61
s38417 46 46 0 0.00 2276 2204 -72 -3.16 1047.1 1070.6 +23.5 +2.25 1.001 1.001 0.000 0.00 1.000 1.003 +0.003 +0.26 0.724 0.717 -0.007 -1.03
s38584.1 21 21 0 0.00 479 471 -8 -1.67 1110.2 1158.2 +48.1 +4.33 1.106 1.106 0.000 0.00 1.082 1.088 +0.006 +0.56 0.737 0.753 +0.016 +2.22
seq 112 111 -1 -0.89 671 560 -111 -16.54 732.0 768.8 +36.8 +5.03 1.130 1.125 -0.005 -0.43 1.126 1.128 +0.002 +0.18 0.596 0.599 +0.003 +0.42
tseng 35 35 0 0.00 636 628 -8 -1.26 478.0 472.2 -5.9 -1.23 1.150 1.151 +0.001 +0.07 1.072 1.107 +0.035 +3.31 0.590 0.577 -0.013 -2.12
Avg. -2.40 -1.65 -199.7 -12.29 -4.69 -1.13 -0.0044 -0.40 +0.0040 +0.37 -0.0027 -0.44
Table 7. Impact of ResynMD on timing cost (total arrival time) and critical-path delay (CPD) across MCNC benchmarks, both reported in ns.
Bench. Timing cost (ns) CPD (ns)
Prev Resysn $\Delta$ $\%$ Prev Resysn $\Delta$ $\%$
apex2 6267.0 6094.5 -173.0 -2.75 42.71 31.16 -11.56 -27.06
apex4 2928.4 2768.9 -159.0 -5.45 42.25 38.42 -3.83 -9.07
bigkey 4285.2 4045.5 -240.0 -5.59 33.75 33.98 +0.23 +0.68
clma 9243.5 10042.0 +799.0 +8.64 34.64 39.41 +4.77 +13.77
des 2896.7 3112.3 +216.0 +7.45 36.37 34.69 -1.67 -4.60
diffeq 1356.4 1183.9 -173.0 -12.72 38.24 31.10 -7.14 -18.68
dsip 2406.0 2386.3 -19.7 -0.82 31.40 27.27 -4.13 -13.15
ex5p 3002.3 2553.5 -449.0 -14.95 46.00 39.55 -6.44 -14.01
frisc 4429.4 4210.9 -219.0 -4.93 33.93 45.84 +11.92 +35.12
misex3 5070.9 5662.7 +592.0 +11.67 36.25 29.44 -6.81 -18.78
pdc 7935.8 7214.6 -721.0 -9.09 43.67 44.07 +0.41 +0.93
s38417 2872.2 2954.1 +81.9 +2.85 29.21 24.77 -4.43 -15.18
s38584.1 2018.8 1852.1 -167.0 -8.25 30.74 30.54 -0.20 -0.64
seq 5001.5 6685.2 +1680.0 +33.66 34.36 38.20 +3.83 +11.15
tseng 918.5 921.8 +3.3 +0.36 28.04 26.98 -1.06 -3.76
Avg. +70.10 +0.01 -1.74 -4.22
Table 8. Impact of ResynMD on physical design across Koios benchmarks with reduced inter-die connectivity (2-die configuration).
Bench. Inter-die connectivity Post-placement metrics
$\Delta$ $N_{\text{sll}}$ $\Delta$ $N_{\text{sll}}^{FO}$ $\Delta$ $BBox_{cost}^{2die}$ (%) $\Delta$ Timing (%) $\Delta$ CPD (%) $\Delta$ Runtime (%)
robot_rl 0 (0.00%) -189 (-10.79%) -2.18 +5.77 $\uparrow$ +0.42 -0.88
tpu_like.medium -2 (-0.61%) -980 (-2.68%) -7.25 $\downarrow$ +3.85 +5.80 $\uparrow$ +0.74
lstm 0 (0.00%) -145 (-0.25%) -8.89 $\downarrow$ -5.50 $\downarrow$ -2.71 -1.35
reduction_layer 0 (0.00%) -8 (-0.10%) -0.17 +4.86 -20.23 $\downarrow$ -2.04
softmax 0 (0.00%) -4 (-0.24%) -2.68 +1.02 -21.42 $\downarrow$ -1.22
dla_like.medium 0 (0.00%) -1 (-0.00%) +1.80 +7.15 $\uparrow$ -1.08 -2.69
dla_like.small 0 (0.00%) -7 (-0.01%) +0.13 +21.92 $\uparrow$ -1.54 -2.70
*Arrows indicate significant variations ($|\Delta|>5\%$).
4.5.1. Inter-Die Connectivity Reduction
Table 6 reports the impact of ResynMD on inter-die connectivity under the 2-die configuration. Across the MCNC (yang1991logic, ) benchmarks, the proposed method reduces inter-die fanout edges by 12.29% on average, while the net-level reduction is 1.65%. The discrepancy between edge-level and net-level improvements is expected. Multiple cross-die sinks may share a single SLL track, and removing individual fanout edges does not necessarily eliminate the occupied SLL channel. Nevertheless, reducing inter-die fanout edges lowers SLL loading and simplifies inter-die connectivity at the structural level.
4.5.2. Placement Quality and Load Balance
The reduction in inter-die connectivity translates to modest improvements in placement quality. On MCNC benchmarks, the average bounding-box cost decreases by 1.13%. Although the average BBox reduction appears modest, it reflects a more spatially compact placement solution. Reducing inter-die fanout edges shortens the effective span of inter-die nets, which decreases routing demand and alleviates congestion in multi-die architectures. Circuits exhibiting larger edge-level reductions (e.g., ex5p and pdc) show more noticeable BBox improvements, suggesting that reduced inter-die coupling can alleviate global placement pressure. These effects are expect to become more pronounced during routing, as reduced congestion could potentially lower the number of rip-up-and-reroute iterations. A detailed routing-level analysis is left for future work.
Die-level load balance remains largely stable. The pre-pack imbalance factor changes by 0.40-0.40% on average, while the post-pack imbalance varies by $0.37\%$. These small variations indicate that interconnect-aware resynthesis does not significantly distort partition balance and remains compatible with partitioning-based multi-die flows. The imbalance change at the LUT-level is smaller than the intrinsic variation introduced by the packing stage. Placement runtime is also stable, with an average change of 0.44-0.44%, demonstrating that the proposed method does not introduce additional overhead in the physical design stage.
In cases of partition imbalance, the optimization can be restricted to nodes in the smaller die subgraph, potentially reducing the effective size of the larger die and improving partition balance. Since the optimization objective adheres to the area constraint defined in Equation 2, the total circuit size is guaranteed to be non-increasing, with only a possible increase in edge count due to additional fanout connections of selected divisors. Therefore, even if the imbalance factor slightly increases (e.g., due to further reduction of the smaller die), the downstream physical design stages are not adversely affected compared with the baseline flow.
4.5.3. Timing Behavior
The impact on timing is circuit-dependent, as shown in Table 7. While several benchmarks (e.g., ex5p, diffeq, and pdc) exhibit timing improvement, some of the others experience degradation. On average, the critical-path delay (CPD) decreases by $4.22\%$. This non-uniform behavior is attributed to the absence of explicit timing constraints in the resynthesis objective. While minimizing inter-die fanout effectively alleviates routing congestion and SLL overhead, logic resubstitution carries the risk of selecting timing-critical signals as divisors for node interpolation. The example in Figure 6(b) illustrates this issue, although the SLL at the pivot node is eliminated, the selected divisor $Y$ lies on a timing-critical path, and its own fanin already crosses die boundaries. This process inadvertently extends the critical path and sabotages timing slack. Therefore, inter-die connectivity reduction alone does not guarantee timing improvement. Incorporating slack-based information as the cost function of SLL-aware logic resynthesis can prevent timing degradation; however, this con will comes at the expense of resubstitution efficacy. These observations highlight the complex interaction between logic-level restructuring and downstream multi-die placement optimization.
4.5.4. Heterogeneous Koios Benchmarks
As shown in Table 8, in Koios benchmarks that contain hard blocks, net-level SLL reductions $\Delta N_{sll}$ are limited, and edge-level improvements $\Delta N_{sll}^{FO}$ are generally small. This is expected, as inter-die connectivity in heterogeneous designs is often dominated by hard macros that cannot be modified by LUT-level resubstitution. In the current Koios flow, the circuits are technology-mapped and timing-optimized prior to resynthesis. Hard blocks and sequential latches are treated as black boxes, and their I/Os are exposed to the LUT-level netlist as primary inputs and outputs. As a result, logical continuity across these modules is structurally interrupted. During window-based resubstitution, such boundaries restrict the availability of valid divisors, since signals cannot propagate across black-box interfaces.
Nevertheless, due to computational resource constraints, we conducted full physical design experiments on 15 Koios circuits, among which 8 exhibit measurable reductions in inter-die connectivity under ResynMD. certain designs such as lstm and tpu_like.medium exhibit BBox reductions of up to 8.89% and 7.25% respectively, with an average of 2.75% reduction on 7 listed circuits. Although net-level SLL reduction is limited in heterogeneous circuits dominated by hard macros, these results indicate that even modest edge-level improvements within combinational regions can reshape spatial placement organization. In terms of timing, the average CPD across the seven listed Koios benchmarks decreases by 5.82%, although the behavior remains design-dependent. Significant improvements are observed in designs such as reduction_layer and softmax, where CPD decreases by over 20%. These results suggest that even when net-level SLL reduction is structurally constrained, localized edge-level simplification can still mitigate inter-die routing overhead and reduce critical-path sensitivity to SLL delay.
These observations imply that SLL-aware resynthesis would be more effective if applied before hard-block binding and physical synthesis. Performing logic restructuring earlier in the flow would provide greater structural freedom and potentially expose more opportunities for reducing inter-die connectivity.
5. Conclusion and Future Work
This paper presents ResynMD, an interconnect-aware logic resynthesis method integrated into a partitioning-based multi-die tool flow to reduce inter-die connections. To the best of our knowledge, this is the first work to incorporate logic resynthesis into a multi-die FPGA CAD flow for inter-die connection optimization. By targeting LUT-level logic that contributes substantially to inter-die connections, ResynMD is orthogonal to existing partitioning-based flows and demonstrates that logic resynthesis can reduce inter-die connections that partitioning alone cannot eliminate. Experimental results show that ResynMD reduces inter-die connections across a wide range of benchmarks, with particularly strong benefits for combinational and soft-logic-dominated designs. These improvements are achieved while maintaining acceptable partition balance and successful physical implementation. The reduction in inter-die connections also leads to downstream effects on physical metrics after placement, such as wirelength and critical path delay.
Future work will incorporate routing into both the evaluation and the resynthesis objective, enabling a more realistic assessment of inter-die routing resource usage and overall implementation quality. Finally, although this work focuses on LUT-based FPGA fabrics, the principle of interconnect-aware logic resynthesis can be extended to chiplet-based ASIC systems or 3D ICs, where inter-die or inter-layer interconnects similarly incur high delay and routing cost. In such settings, the optimization would operate on technology-mapped gate-level netlist rather than on LUT-level netlists, by performing functionality-preserving transformations to reduce inter-die or inter-layer connections.